OpenStudy (anonymous):

solve the proportion- 16/50=x/156.25

OpenStudy (anonymous):

\[\huge \frac{16}{50}= \frac{x}{156.25}\]

OpenStudy (anonymous):

\[\huge \frac{16}{50}= \frac{x}{156.25}\] \[\huge \frac{16}{50}= \frac{x}{156.25} \rightarrow x= \frac { 16 \times 156.25}{50}\] \[\huge \rightarrow x= 16 \times 3.125 = 50\]

OpenStudy (anonymous):

@js777

OpenStudy (anonymous):

wow! can you put it in simple words lol @dpasingh

OpenStudy (anonymous):

\[x=(16\times156.25)/50\]

OpenStudy (anonymous):

thanks @Shirin74

OpenStudy (johnweldon1993):

You have the proportion... \[\large \frac{16}{50} = \frac{x}{156.25}\] You cross multiply...do 16 times 156.25 and then 50 times x Now you have \[\large 2500 = 50x\] Just divide both sides by 50 to get \[\large x = 50\]

OpenStudy (anonymous):

thanks @johnweldon1993

OpenStudy (anonymous):

Sure... Carry out the cross multiplication , so it would be 50*x = 16* 156.25 now divide both sides by 50 to find x= 16* 156.25 / 50 next divide 156.25/ 50 we get 3.125 multiply 16 by 3.125 to get 50 as answer hence x= 50

OpenStudy (anonymous):

thanks! @dpasingh

OpenStudy (anonymous):

lol kk @dpasingh

OpenStudy (anonymous):

kk lol @dpasingh

OpenStudy (anonymous):

your welcome honey @js777